# Towards Passive Safe Reinforcement Learning: A Comparative Study on Contact-rich Robotic Manipulation

### v1

@@ III-C Passivity-based approach for variable stiffness control
$\dot{S}_{\text{Ctrl}}+P_{\text{IC}}=0.$ (3)
@@
$\dot{S}_{\text{Ctrl}}+\dot{E_{\text{T}}}+P_{\text{D}}\leq 0,$ (4)
@@
In sections III-D and III-E, we go into more detail on how we incorporate tank and flow conditions into the learning and deployment phases, respectively.
III-D Passivity-aware safe RL
This section describes how we add passivity awareness to the RL policy. As described in Section III-A, the safe RL framework achieves safety through constraint definitions. We extend the CMDP formulation by including the passivity constraints so that the safety critic also learns the concept of passivity. As illustrated in Fig. 1.b, this information is provided as the energy constraints ${C_{e}}\subset\mathcal{C}$, including a budget constraint ${c_{eb}\in{C_{e}}}$ and a flow constraint ${c_{ef}\in{C_{e}}}$, which correspond to the tank and flow conditions.The budget constraint ${c_{eb}}$ is introduced to make the agent aware of the amount of energy it has available for adapting the Cartesian stiffness $\mathbf{K}$ during the task. The constraint is triggered whenever $E_{\text{T}}+\dot{E_{\text{T}}}\leq\underline{E}_{\text{T}}$, which indicates that the lower bound of the energy tank is reached.
Whereas the purpose of introducing the energy flow constraint ${c_{ef}}$ is to make the agent aware of how much energy it is allowed to take out of the tank at any given time for a specific task. Therefore ${c_{ef}}$ is triggered whenever the draining energy flow of the energy tank exceeds the defined threshold $\dot{E}_{\text{T}}<\underline{\dot{E}}_{\text{T}}$. RL training episodes are terminated when a constraint happens. This makes it easy to end an episode with both of the energy constraints, hindering the training process. Thus, we experiment with different combinations of these constraints as detailed in Sec. IV. Defining the passivity information as constraints means that the training data includes this information. This data is used to train the safety critic, which learns the risk related to passivity. Therefore, we expect the passivity-aware RL agent to learn to use the energy more economically, in means of cumulative energy spent and instant energy flow. However, learning a regression model does not guarantee that the risk is always predicted correctly. Thus, we recommend deploying these models with the added passivity layer for passivity-ensured control, as described in the following section.
III-E Passivity-ensured control
We employ a passivity layer at the output of the RL agent to guarantee the control passivity, as illustrated in Fig. 1.c. The passivity layer contains two filtering actions to satisfy the tank and flow conditions defined in Section III-C. We ensure the flow condition [13] by limiting the rate at which the controller can inject energy into the system through the introduction of the scaling variable
$\alpha=\begin{cases}\frac{\underline{\dot{E}}_{\text{T}}}{\dot{E}_{T}}&\text{if }\dot{E}_{T}<\underline{\dot{E}}_{\text{T}}\leq 0\\ 1&\text{otherwise}.\end{cases}$ (5)
@@ III-E Passivity-ensured control
$\dot{E_{\text{T}}}^{\prime}=\begin{cases}\alpha\dot{E_{\text{T}}}&\text{if }\underline{E}_{\text{T}}\leq E_{\text{T}}+\dot{E_{\text{T}}}\leq\overline{E}_{\text{T}}\\ 0&\text{otherwise}.\end{cases}$ (6)
Integrating the described energy tank dynamic for the VIC, the change of the Cartesian stiffness can be defined as
$\mathbf{K}(t)=\begin{cases}\mathbf{K}(t{-}1)&\text{if }E_{\text{T}}+\dot{E_{\text{T}}}^{\prime}\leq\underline{E}_{\text{T}}\\ \mathbf{K}(t)&\text{otherwise}.\\ \end{cases}$ (7)
@@
${\boldsymbol{\tau}}^{\top}_{\text{IC}}=\alpha{\mathbf{J}}^{\top}(\mathbf{q})\big{(}{\mathbf{w}^{0,\text{EE}}_{K}}^{\top}+\mathbf{D}\dot{\mathbf{x}}\big{)}.$ (8)
@@
IV Evaluation
We evaluate our method through the contact-rich maze exploration task. In this task, the robot has to explore an unknown maze blindly. Thus, it relies on the force sensing for finding its way out, making it a highly contact-rich task, i.e., the robot cannot avoid contact and it has to regulate the exerted wrenches for safety. In the following evaluation experiments, we aim to study: 1. the passivity from an energy tank perspective during policy learning; 2. how passivity constraints affect RL policy training and performance; 3. how to guarantee both safe and stable performance while executing the task efficiently.
We conduct three types of experiments to study these points. First, we train the agent with different energy tank and energy flow constraints, and compare them to the baseline training that is agnostic to passivity (Sec. IV-B). Second, we deploy the learned models with the addition of the passivity layer in the control loop as shown in Fig. 1 (Sec. IV-C). The passivity layer guarantees control stability by limiting the actions of the agent. Our goal is to understand the trade-off between safety and performance in scenarios that are aware or agnostic of the passivity constraints. Third, we deploy the passivity-integrated approach in a real-world contact-rich robotic task to validate the applicability of our approach (Sec. IV-D).
In the following text, we first describe our experiment setup, then we present the details and results of these three experiments in the three subsequent sections.
@@ IV Evaluation
IV-A Experiment setup
The experimental task is maze exploration through interaction, where the end-effector trajectories and stiffness values adapt dynamically. We train the policy in simulation using a Mujoco (version 2.3.3) maze environment, adapted from previous work [6], as shown in Fig. 2. The objective is to navigate through the maze only by relying on contact forces, moving from the start point on the left to the goal point on the right. During this process, the controller dynamically adjusts Cartesian stiffness and pose in response to disturbances, such as maze walls and movable obstacles represented by three simulated sphere balls in the middle of the maze. We train the policy on the cluster (1×1$\times$1 × NVIDIA Tesla V100 16Gb GPU, 1×1$\times$1 × Intel(R) Xeon(R) Silver 4210 CPU)  [^1] We gratefully acknowledge the Data Science and Computation Facility and its Support Team at Fondazione Istituto Italiano di Tecnologia.. For the real-world experiment setup please refer to IV-D.
TABLE I: Experiment parameters
@@ IV-A Experiment setup
Figure 3: Passivity-aware policy training experiment results. We present the total numbers of success and violation in the first two plots. The third plot shows the ratio of mean success to mean violation. We zoom into the last 5 case results as they are almost invisible.
IV-B Passivity-aware policy training
[…]
We also train the passivity-agnostic approach as the baseline (agnostic) without any energy-based constraints but only with contact-force constraint as shown in Fig. 1, thus the task is more easily achieved during training, however, the policy stays unaware of system stability.
IV-B1 Offline data collection
We collected 40,000 tuples using the same method as in [6], each consisting of [$s$, $a$, $s_{t+1}$, mask] as offline data, where mask indicates a constraint violation. In total, there are 1363 constraint violations. In comparison to [6], we increased the force threshold to 40 N (same as in  [26]) in this study to better focus on the passivity-based constraints.
IV-B2 Online training
We first train a passivity-agnostic policy where the agent focuses on safe task completion but is agnostic to system stability. For comparing it to the passivity-aware policies where the budget and energy flow constraints are added during training. These energy constraints, along with the contact force constraint, constitute three distinct constraints as part of CMDP. We trained policies through SAC [27] (see the basic hyper parameters in Tab. I) under different constraint configurations. To minimize random error and statistical bias, we train each policy under various constraint configurations using 10 different random seeds.
IV-B3 Results
The results of the training experiments are summarized as means and standard deviations of 10 seeds in Fig. 3. As the most relaxed configuration, agnostic solves the task earlier than others. We see a consistent increase in constraint violations as the budget size decreases and harsher flow limits are applied. Flow constraints dominate the type of violations as in Eb8 against Eb8-Ef7. These findings suggest that the passivity-constrained cases would benefit from further training.
@@ IV-B3 Results
Figure 4: These plots visualize the effect that the changes of the stiffnesses $k_{T_{x}},k_{T_{y}}\in\mathbf{K}$ during training, have on the energy flow $\dot{E_{\text{T}}}$ and consequently on the energy stored in the tank $E_{\text{T}}$.
During training as well as deployment, the agent regulates the stiffness of the impedance controller (Sec. III-B). This has a direct effect on how the energy flow $\dot{E_{\text{T}}}$ and the tank budget $E_{\text{T}}$, which can be seen in Fig. 4. More specifically, it can be seen that whenever the agent increases the stiffness of springs during its motion, it results in $\dot{E_{\text{T}}}<0$ as long as $\underline{E}_{\text{T}}\leq E_{\text{T}}<\overline{E}_{\text{T}}$ holds as described in Sec. III-C. Furthermore, it can be observed that the greater the increase in stiffness during a particular motion the greater effect it has on $\dot{E_{\text{T}}}$ and consequently on $E_{\text{T}}$. This behavior can be directly related to the definition of $\dot{E_{\text{T}}}$ (Sec. III-C). During training, it can observed how the agent becomes more aware of the effect the changes of $\mathbf{K}$ has on the either $\dot{E_{\text{T}}}$ and/or $E_{\text{T}}$ depending on the constraints it is subjected to.
@@
Figure 5: Plots illustrate the effect that the introduction of the tank constraint has on the agents energy expenditure at different stages during training.
@@
We present the impact of energy constraints in passivity-aware policy training in Fig. 5, showing the agent’s energy expenditure at two different stages during training. Specifically, the red graph is the energy expenditure and the energy flow at the port of the tank in the first episode, in other words it has not learned to adapt to this constraint. The blue graph represents the agent’s energy-related behavior after 100 episodes, making it evident that the agent has learned to act more conservatively, resulting in reduced and more evenly distributed energy expenditure. Hence, it can be seen how introducing the energy constraint has a direct impact on the agent’s behavior during training.
After training, we test our policy on representative experiments, including agnostic, Eb4, Eb8, Eb8-Ef7, for comparison. First, all policies are evaluated over 100 episodes under the same constraint settings used during training. The testing results, shown in Table II, demonstrate the model’s strong performance under conditions consistent with its training environment. Notably, the agent achieves a 100% success rate when it is agnostic to energy constraints, as it is not subjected to $C_{e}$ and can therefore operate more exploitatively during its runs. Furthermore, a comparison between Eb8 and Eb4 reveals that the more conservative tank size in Eb4 consequentiality results in more tank violations. Additionally, it can be seen that the combination of both tank and flow constraints (Eb8-Ef7) results in overall worse performance.
@@
TABLE II: Results of model tests. Performance evaluation over 100 episodes for each of three randomly selected seeds.
@@
Figure 7: Energy budget progress during passivity-ensured deployment experiments. The curves indicate the mean of 300 runs (100 per seed) and the shades indicate the standard deviation. Although the passivity-agnostic agent is kept stable by the passivity layer, it shows no awareness of energy expenditure.
@@
Figure 8: Snapshots in Real-world Experiments.(a) the overview of maze exploration task, at start point. (b) end of task. (c) closeup robot moves along with contact force and executes at the turnaround area. (d) failure case. violation at turnaround areas, where the robot tries to keep stable and safe but too conservatively.
IV-C Passivity-ensured deployment
@@
In the deployment experiments, we evaluate the effect of the passivity layer on the trained models, by comparing the performances of the passivity-agnostic baseline and various passivity-aware approaches in combination with and without the passivity layer. In these experiments, we do not retrain the models but impose the passivity layer on the output of the RL agent (Fig. 1), where the layer is with uniform parameter settings: an initial energy budget of $\overline{E}_{\text{T}}{=}6$ and an energy flow limit of $\dot{\underline{E}_{\text{T}}}{=}-0.5$.
@@
The results in Fig. 6 showcase the task successes, and violations. Please note that violations here only indicate force violations. No energy constraint violation happens because of the passivity layer. Overall, the results labeled as “with passivity layer” outperform the baseline labeled as “no passivity layer” in Fig. 6. These results reveal the risks of deploying models without passivity guarantees, leading to system instability and unsafe behaviors despite good performance in model tests shown in Table II. Specifically: i) The agnostic policy performs the worst when no passivity filter is applied, as it consistently violates the flow constraint. Although it achieves a 100% success rate when evaluated without considering energy constraints shown in Table II, which highlights the necessity of incorporating passivity during training. ii) The success rate improves significantly when the passivity filter is applied underscoring the dual benefits of passive-aware policies and the passivity layer, ensuring safer and more stable deployments.
@@
We also present the energy usage of each method in Fig. 7 to further demonstrate the benefit of passivity-aware training. Even though all of agnostic, eb8 and eb4 show full success in the tests (Fig. 6), analysis of the energy budget progress in these runs reveals that the agnostic model depletes its budget rapidly. This rapid depletion reduces its likelihood of success in more complex task instances. On the other hand, eb4 demonstrates the best success and energy efficiency, despite its delayed learning shown in Fig. 3.
IV-D Real-world experiments

### v2

@@ II-C Passivity-based approach for variable stiffness control
$\dot{S}_{\text{Ctrl}}+P_{\text{IC}}=0.\vspace{-0.1cm}$ (3)
@@
$\dot{S}_{\text{Ctrl}}+\dot{E_{\text{T}}}+P_{\text{D}}\leq 0,\vspace{-0.1cm}$ (4)
@@
In sections II-D and II-E, we go into more detail on how we incorporate energy tank and energy flow conditions into the learning and deployment phases, respectively.
II-D Passivity-aware safe RL
This section describes how we add passivity awareness to the RL policy. As described in Section II-A, the safe RL framework achieves safety through constraint definitions. We extend the CMDP formulation by including the passivity constraints so that the safety critic also learns the concept of passivity. As illustrated in Fig. 1.b, this information is provided as the energy constraints ${C_{e}}\subset\mathcal{C}$, including a budget constraint ${c_{eb}\in{C_{e}}}$ and a flow constraint ${c_{ef}\in{C_{e}}}$, which correspond to the energy tank and energy flow conditions.The budget constraint ${c_{eb}}$ is introduced to make the agent aware of the amount of energy it has available for adapting the Cartesian stiffness $\mathbf{K}$ during the task. The constraint is triggered whenever $E_{\text{T}}+\dot{E_{\text{T}}}\leq\underline{E}_{\text{T}}$, which indicates that the lower bound of the energy tank is reached.
Whereas the purpose of introducing the energy flow constraint ${c_{ef}}$ is to make the agent aware of how much energy it is allowed to take out of the energy tank at any given time for a specific task. Therefore ${c_{ef}}$ is triggered whenever the draining energy flow of the energy tank exceeds the defined threshold $\dot{E}_{\text{T}}<\underline{\dot{E}}_{\text{T}}$. RL training episodes are terminated when a constraint happens. This makes it easy to end an episode with both of the energy constraints, hindering the training process. Thus, we experiment with different combinations of these constraints as detailed in Sec. III. Defining the passivity information as constraints means that the training data includes this information. This data is used to train the safety critic, which learns the risk related to passivity. Therefore, we expect the passivity-aware RL agent to learn to use the energy more economically, in means of cumulative energy spent and instant energy flow. However, learning a regression model does not guarantee that the risk is always predicted correctly. Thus, we recommend deploying these models with the added passivity layer for passivity-ensured control, as described in the following section.
II-E Passivity-ensured control
We employ a passivity layer at the output of the RL agent to guarantee the control passivity, as illustrated in Fig. 1.c. The passivity layer contains two filtering actions to satisfy the energy budget and energy flow conditions defined in Section II-C. We ensure the flow condition [8] by limiting the rate at which the controller can inject energy into the system through the introduction of the scaling variable
$\alpha=\begin{cases}\frac{\underline{\dot{E}}_{\text{T}}}{\dot{E}_{T}}&\text{if }\dot{E}_{T}<\underline{\dot{E}}_{\text{T}}\leq 0\\ 1&\text{otherwise}.\vspace{-0.1cm}\end{cases}$ (5)
@@ II-E Passivity-ensured control
$\dot{E_{\text{T}}}^{\prime}=\begin{cases}\alpha\dot{E_{\text{T}}}&\text{if }\underline{E}_{\text{T}}\leq E_{\text{T}}+\dot{E_{\text{T}}}\leq\overline{E}_{\text{T}}\\ 0&\text{otherwise}.\end{cases}\vspace{-0.1cm}$ (6)
Integrating the described energy tank dynamics for the VIC, the change of the Cartesian stiffness can be defined as
$\mathbf{K}(t)=\begin{cases}\mathbf{K}(t{-}1)&\text{if }E_{\text{T}}+\dot{E_{\text{T}}}^{\prime}\leq\underline{E}_{\text{T}}\\ \mathbf{K}(t)&\text{otherwise}.\\ \end{cases}\vspace{-0.1cm}$ (7)
@@
${\boldsymbol{\tau}}^{\top}_{\text{IC}}=\alpha{\mathbf{J}}^{\top}(\mathbf{q})\big{(}{\mathbf{w}^{0,\text{EE}}_{K}}^{\top}+\mathbf{D}\dot{\mathbf{x}}\big{)}.\vspace{-0.1cm}$ (8)
@@
III Evaluation
We evaluate our method through the contact-rich maze exploration task. In the following evaluation experiments, we aim to study: 1. how passivity constraints affect RL policy training and performance; 2. how to guarantee both safe and stable performance while executing the task efficiently. To do so, we conduct three types of experiments. First, we train the agent with different energy budget and energy flow constraints, comparing them to the passivity-agnostic baseline (Sec. III-B). Second, we deploy the learned models with the passivity layer (Sec. III-C), which guarantees control stability by limiting agent actions. Third, we validate our approach through real-world experiments (Sec. III-D).
@@ III Evaluation
III-A Experimental design
The experimental task is a maze exploration task where the robot has to explore an unknown maze blindly. Thus, it relies on the force sensing for finding its way out, making it a contact-rich task, i.e., the robot does not avoid the contact and it has to regulate the exerted wrenches for safety. During this process, the controller dynamically adjusts Cartesian stiffness and pose in response to disturbances, such as maze walls and movable obstacles represented by three simulated sphere balls in the middle of the maze.
[…]
TABLE II: Experiment parameters
@@ III-A Experimental design
Figure 3: Passivity-aware policy training experiment results. We present the total numbers of success and violation in the first two plots. The third plot shows the ratio of mean success to mean violation.
III-B Passivity-aware policy training
As mentioned previously, this section focuses on how the passivity constraints effect the training of a RL policy, by comparing them to each other as well as to the passivity-agnostic approach that is solely subjected to the contact-force constraint as shown in Fig. 1. More specifically the remainder will focus on energy budgets that are within the range of $4{-}10$ and flow limits are from $-0.7$ to $-0.3$, which were found based on preliminary trials. The results of the training experiments are summarized as means and standard deviations of 10 seeds in Fig. 3. More specifically, it shows the amount of success in terms of task completion, the type and number of constraint violations which let to the task termination as well as the ratio between the amount of successfully completed runs and violations.
Firstly, it can be seen that the tasks success is directly correlated with the amount of constraints introduced to the CMDP. Secondly, the introduction of $c_{eb}$ has a direct effect on the amount of force violation, which overall decreases as the energy budget gets smaller. Furthermore, the introduction of $c_{ef}$ results in that the force constraint are no longer violated. The only outlier is the Eb10, which outperforms the agnostic run, even though it violates both the budget and force constraint. Thirdly, it can be seen that the introduction of $c_{ef}$ has a more aggressive effect on the task success than the introduction of only $c_{eb}$. This is due to the fact that the constraint on the flow hinders the agent to make aggressive changes in its stiffnesses, resulting in an overly conservative approach which hinders the agent to explore the maze. Hence, generally it can be said that there is a consistent increase in constraint violations as the budget size decreases and harsher flow limits are applied.
@@ III-B Passivity-aware policy training
Figure 4: These plots visualize the effect that the changes of the stiffnesses $k_{t_{x}},k_{t_{y}}\in\mathbf{K}_{\text{T}}$ during training, have on the energy flow $\dot{E_{\text{T}}}$ and consequently on the energy stored in the energy tank $E_{\text{T}}$.
@@
Figure 5: Plots illustrate the effect that the introduction of the energy tank constraint has on the agents energy expenditure at different stages during training.
During training the agent regulates the stiffness of the impedance controller (Sec. II-B). This has a direct effect on how the energy flow $\dot{E_{\text{T}}}$ and the energy tank budget $E_{\text{T}}$, which is visualized in Fig. 4. More specifically, it is shown that whenever the agent increases the stiffness of springs during its motion, it results in $\dot{E_{\text{T}}}<0$ as long as $\underline{E}_{\text{T}}\leq E_{\text{T}}<\overline{E}_{\text{T}}$ holds as described in Sec. II-C.
@@
Furthermore, it can be observed that the greater the increase in stiffness during a particular motion the greater effect it has on $\dot{E_{\text{T}}}$ and consequently on $E_{\text{T}}$. This behavior can be directly related to the definition of $\dot{E_{\text{T}}}$ (Sec. II-C). During training, one can observe how the agent becomes more aware of how the change of $\mathbf{K}$ effects either $\dot{E_{\text{T}}}$ and/or $E_{\text{T}}$ depending on the constraints it is subjected to. The impact that $C_{e}$ has on passivity-aware policy training can be seen in Fig. 5, showing the agent’s energy expenditure at two different stages during training. Where the red graph representing the energy expenditure and the energy flow at the port of the energy tank in the first episode, and the blue graph represents the agent’s energy-related behavior after 100 episodes. When comparing both graphs it is evident that the agent has learned across these 100 episodes to act more conservatively, resulting in reduced and more evenly distributed energy expenditure. Hence, it can be seen how introducing the energy constraints $C_{e}$ have a direct impact on the agent’s behavior during training.
@@
TABLE III: Results of model tests. Performance evaluation over 100 episodes for each of three randomly selected seeds.
@@
After training, we selected the policies for agnostic, Eb4, Eb8, Eb8-Ef7 to test the model in representative experiments. All policies are evaluated over 100 episodes under the same constraint settings used during training. In the testing results, shown in Table III, similar to the training results shown in Figure 3, the agent achieves the highest success rate when it is agnostic to $C_{e}$ and can therefore operate more exploratory during its runs. Furthermore, looking at Eb8 and Eb4, a more conservative energy tank size results in more tank violations. However, compared to the training results, the Eb8-Ef7 run obtained less successes as it has harsher constraints during results. Here, Eb4 obtains high success, even though it had fewer successes than Eb8-Ef7 in training. From this it can be concluded that Eb4 learned a better task policy, although it was delayed by its limited energy budget.
@@
[…]
Figure 7: (a) Energy budget progress during passivity-ensured deployment experiments. The curves indicate the mean of 300 runs (100 per seed) and the shades indicate the standard deviation. (b) Histogram of the run times, i.e., the number of runs that finished in time duration intervals.
[…]
@@
Figure 8: Snapshots and result analysis in real-world experiments. (a) the overview of maze exploration task. (b) closeup when robot moves along with contact force. (c) failure case agnostic but non-passive behavior at the turnaround area. (d) shows the energy budget consumption among different setups, demonstrating our method outperforms in terms of passivity and energy efficiency.
III-C Passivity-ensured deployment
@@
In the deployment experiments, we evaluate the effect of the passivity layer on the trained models, by comparing the performances of the passivity-agnostic baseline and various passivity-aware approaches in combination with and without the passivity layer. Where the passivity-agnostic approach with the passivity layer is called passivity-filtered and the passivity-aware approaches with the filter become passivity-ensured. In these experiments, we do not retrain the models but impose the passivity layer on the output of the RL agent (Fig. 1), where the layer is with uniform parameter settings: an initial energy budget of $\overline{E}_{\text{T}}{=}6$ and an energy flow limit of $\dot{\underline{E}_{\text{T}}}{=}-0.5$.
The results in Fig. 6 showcase the task successes and violations. Overall, the results labeled as “with passivity layer” outperform the baseline labeled as “no passivity layer”. These results reveal the risks of deploying models without passivity guarantees, which potentially leads to system instabilities and unsafe behaviors despite good performance in model tests shown in Table III.
@@
Specifically: 1. The agnostic policy performs the worst when no passivity filter is applied, as it consistently violates the flow constraint. Although it achieves a 100% success rate when evaluated without considering energy constraints shown in Table III, which highlights the necessity of incorporating passivity during training. 2. The success rate improves significantly when the passivity filter is applied, underscoring the dual benefits of passive-aware policies and the passivity layer, ensuring safer and more stable deployments. 3. Violations dropped significantly with the passivity layer, with only Eb8 had two force violations that is almost not visible in Fig. 6. Note that no energy constraint violations occured, as the passivity layer ensures system passivity.
@@
We also present the energy usage of each method in Fig. 7.a to further demonstrate the benefit of passivity-aware training. Even though all of agnostic, Eb8 and Eb4 show full success in the tests (Fig. 6), analysis of the energy budget progress in these runs reveals that the agnostic model depletes its budget more rapidly. This rapid depletion reduces its likelihood of success in more complex task instances. On the other hand, Eb4 demonstrates the best success and energy efficiency, despite its delayed learning shown in Fig. 3. Furthermore, the histogram in Fig. 7.b shows number of runs distribution on completion time. It is clear that both Eb8 and agnostic agents finish the task around 30s in majority of runs, while some runs take longer due to getting stuck in the maze. This highlights that conservative agents (Eb4, Eb8-Eb7) take slightly longer time to finish the maze on average.
III-D Real-world experiments